OpenStudy (anonymous):

For which pair of launch angles will two identical projectiles have equal ranges? A. 19.24°, 80.54°B. 16.42°, 74.58°C. 60.23°, 29.77°D. 89.53°, 01.47°E. 42.42°, 47.59°

OpenStudy (anonymous):

If we establish the kinematic equations of motion, we see that\[y(t) = v_ot \sin(\theta) - {1 \over 2} g t^2\]and\[x(t) = v_o t \cos(\theta)\] Setting \(y(t) = 0\) yields \(0=v_ot \sin(\theta) - {1 \over 2} g t^2\). If we solve for \(t\), we obtain, by factoring, \[t = {2 v \sin(\theta) \over g}\]. Substitute this into our equation for \(x(t)\). This yields\[x(t) = {2 v^2 \cos(\theta) \sin(\theta) \over g}\]Recall that \[\sin(2\theta) = 2\sin(\theta) \cos(\theta)\]Therefore, \[x = {v^2 \sin(2\theta) \over g}\]Assuming the same launch velocity and knowing that gravity is constant, we can establish the two angles from \(\sin(2 \theta)\)

OpenStudy (anonymous):

I'm am so confused... So I multiply sin 20 by the angles?

OpenStudy (calculusfunctions):

He's right. In simple terms multiply each value in each choice by two, and then find their respective sine ratios. See if they're equal.

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

still dont understand uggh XO